At a business college, 80 students are selected ar random. Of the 20 female, 5 will major in Statistics, while of the 60 males, 15 will major in Statistics. If one of the 80 students is selected at random, find the probability that the students selected will:
Not major in Statistics and not be a male

Answer by stanbon(75887)

At a business college, 80 students are selected ar random. Of the 20 female, 5 will major in Statistics, while of the 60 males, 15 will major in Statistics. If one of the 80 students is selected at random, find the probability that the students selected will:
Not major in Statistics and not be a male
---------
Not male = 20 females
Not male and not major in stat 20-5 = 15
----
P(not male and not stat) = 15/80 = 3/16
-----
Cheers,
Stan H.
--------------